OpenStudy (anonymous):

Solve for x in the proportion x+1 over 8 and 2x - 5 over 2

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {x + 1 \over 8} = {2x - 5 \over 2} }\) Let's now cross multiply to eliminate the fraction. \(\Large \color{MidnightBlue}{\Rightarrow 2(x + 1) = 8(2x - 5) }\) \(\Large \color{MidnightBlue}{\Rightarrow 2x + 2 = 16x - 40 }\) \(\Large \color{MidnightBlue}{\Rightarrow 2x + 42 = 16x }\) \(\Large \color{MidnightBlue}{\Rightarrow 14x = 42 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 3 }\)

Parth (parthkohli):

@Brad1996 Do you know why we cross multiply?

OpenStudy (anonymous):

i forgot the reason i just know ur suppose to

OpenStudy (anonymous):

@ParthKohli : suppose u multiple equation by multiple product of each side dem . i mean for this ex suppose u multiple equation by (2*8) u cancel num and dem and it;s just like u cross multiple .

Parth (parthkohli):

Of course @Roya I know that. I was asking Brad.

OpenStudy (anonymous):

@ParthKohli , sri dear .

Parth (parthkohli):

He he no need